OpenStudy (anonymous):

Calculate the center of mass of the area between f(x)=x and f(x)=x^2 mass=12x

OpenStudy (anonymous):

i dont have time to answer but others may find the diagram useful 

OpenStudy (anonymous):

i know the diagram, im just confused of calculating the Center of Mass Coordinates and the moments of y and x.

OpenStudy (dumbcow):

isn't the center of mass found by this sum: \[\int\limits_{0}^{1}x*f(x) dx\] where f(x) is height of region f(x) = x-x^2

OpenStudy (anonymous):

i haven't actually done this kind before but drawing from what i know of moments would it be: \[\frac{\int\limits{12x(x-x^2)x}dx}{\int\limits{12x(x-x^2)}dx}\]

OpenStudy (anonymous):

aah i have to go

OpenStudy (anonymous):

the mass dm=(12x)(x-x^2)dx

OpenStudy (anonymous):

aw ok

OpenStudy (anonymous):

moment of y=∫1/2(x^2)(12x)(x-x^2)dx from 0 to 1 right?

OpenStudy (anonymous):

that would be moment of x, my bad. so the solution would give me y bar. am i correct?

OpenStudy (anonymous):

let (x',y') be the centroid. then \[x'=\ M_y/Area\] \[y'=\ M_x/2/Area\] where M_y is moment about y axis and M_x is moment about x axis find first Area bounded by the graphs: \[A=\int\limits_{0}^{1}x-x^2dx=(x^2/2 -x^3/3)|_{0}^{1}=1/6\] Moment about the x axis: \[1/2\int\limits_{0}^{1}x^2-x^4dx=1/2(x^3/3-x^5/5)|_{0}^{1}=1/2(2/15)=1/15\] Moment about y-axis: \[\int\limits_{0}^{1}(x-x^2)xdx=\int\limits_{0}^{1}x^2-x^3dx=(x^3/3-x^4/4)|_{0}^{1}=1/12\] x'=1/12/1/6=1/2 y'=1/15/1/6=6/15 so our centroid is (1/2,6/15)

OpenStudy (dumbcow):

here is a good reference http://tutorial.math.lamar.edu/Classes/CalcII/CenterOfMass.aspx

OpenStudy (anonymous):

@anonymoustwo44 but you didnt even use the rate of mass. it wasnt constant it was 12x

OpenStudy (anonymous):

@Zarkon how would you factor in the rate of density?

OpenStudy (anonymous):

wouldn't the mass be \[m = \int\limits{(x - x^2) \times 12x}dx\]